the difference between two number is 834. If the larger number is 4 times as much as the smaller number, what is the sum of the two numbers?
Answer by ilana(307)

I'm going to use the variables x and y with x larger than y. So My 2 equations are x - y = 834 and x = 4y. Using the substitution method, I get 4y - y = 834, so 3y = 834, so y = 278. Then x = 4*278, so x = 1112. We want x+y, so x+y = 278 + 1112 = 1390.